OpenStudy (anonymous):

Please help me solve this question.State two ways in which a dimensionally consistent equation may be physically incorrect.

OpenStudy (vishweshshrimali5):

Kinetic energy = mv^2

OpenStudy (anonymous):

???

OpenStudy (vishweshshrimali5):

See the dimensions match on LHS and RHS but we know that equation is incorrect.

OpenStudy (vishweshshrimali5):

Similarly, F = mv^2/r^2 S = ut - at^2

OpenStudy (anonymous):

ok...i understand...but i think u didn't answer my question...

OpenStudy (vishweshshrimali5):

which part ?

OpenStudy (anonymous):

the top one...

OpenStudy (vishweshshrimali5):

top one ?

OpenStudy (vishweshshrimali5):

Sorry I can't get you.

OpenStudy (anonymous):

ok...nevermine...but thanks for ur explain...

OpenStudy (vishweshshrimali5):

Well it was my pleasure. Never hesitate to ask if you couldn't understand any part of the solution. :) Good day

OpenStudy (anonymous):

ok...Thanks...